# Length-Induced Embedding Collapse in Transformer-based Models

### v1

@@ B.3 Proof of Corollary 4
where $\alpha_{1}$ and $\alpha_{2}$ represent the maximum values in the frequency domain of $\mathcal{HC}\left[{\bm{x}_{1}}\right]$ and $\mathcal{HC}\left[{\bm{x}_{2}}\right]$ and $\alpha$ represent the value of $\mathcal{DC}\left[{\bm{x}_{1}}\right]$ and $\mathcal{DC}\left[{\bm{x}_{2}}\right]$, respectively, after applying the discrete Fourier transform. Eqn. 24 leverages that $\mathcal{HC}\left[{\cdot}\right]$ and $\mathcal{DC}\left[{\cdot}\right]$ are orthogonal. Eqn. 25 leverages that the assumption the mean of word embeddings in natural language texts maintains a relatively consistent representation. Thus $\mathcal{HC}\left[{\bm{x}_{1}}\right]$ and $\mathcal{HC}\left[{\bm{x}_{2}}\right]$ have the same value $\alpha$ after applying the discrete Fourier transform. Finally, according to Theorem 3, $\alpha_{1}$ and $\alpha_{2}$ will gradually decrease with $n$ grows, leading to a higher cosine similarity between $\bm{x}_{1}$ and $\bm{x}_{2}$.
Appendix C Datasets and Evaluation Metrics
@@
C.1 Classification
AmazonPolarity (Zhang et al., 2015) consists of Amazon customer reviews, each labeled as either “positive" or “negative."
Banking77  (Casanueva et al., 2020) dataset consists of online banking user queries labeled with one of 77 specific intents.
Emotion  (Saravia et al., 2018) comprises Twitter messages categorized by six fundamental emotions: anger, fear, joy, love, sadness, and surprise.
Imdb  (Maas et al., 2011) consists of extensive movie reviews categorized as either positive or negative.
MassiveIntent  (FitzGerald et al., 2022) is a multilingual dataset featuring a diverse array of utterances from Amazon Alexa, each labeled with one of 60 different intents across 51 languages.
MassiveScenario  (FitzGerald et al., 2022) dataset comprises a diverse collection of Amazon Alexa user utterances, each labeled with one of 60 thematic intents, and supports 51 languages.
TweetSentimentExtraction  [^2] TweetSentimentExtraction, a dataset from a Kaggle competition focuses on classifying tweets into three categories: neutral, positive, and negative sentiments.
C.2 Clustering
Clustering aims at grouping a given set of sentences or textbfs into meaningful clusters by training a mini-batch k-means model on the text embeddings. The model is scored using the v-measure  (Rosenberg & Hirschberg, 2007). Since the v-measure does not depend on the cluster labels, the arrangement of the labels will not affect the score.
ArxivClusteringS2S, ArxivClusteringP2P, BiorxivClusteringS2S, BiorxivClusteringP2P, MedrxivClusteringP2P, MedrxivClusteringS2S (Muennighoff et al., 2023). These datasets are tailored for MTEB, utilizing titles or a combination of titles and abstracts from arXiv, bioRxiv, and medRxiv, with clustering labels derived from human-assigned categories, emphasizing both main and secondary classification levels.
RedditClustering (Geigle et al., 2021), a dataset consists of titles from 199 subreddits, is organized into 25 splits, each featuring 10 to 50 classes, with every class containing between 100 and 1,000 sentences.
RedditClusteringP2P (Muennighoff et al., 2023), developed for the MTEB, consists of Reddit posts combined with their titles, organized into ten splits featuring 10 and 100 clusters each, with a total of 1,000 to 100,000 posts, aimed at clustering based on subreddit affiliation.
StackExchangeClustering (Geigle et al., 2021), a dataset consisting of titles from 121 Stack Exchange communities, is organized into 25 subsets, each containing 10 to 50 categories, with 100 to 1,000 sentences per category.
StackExchangeClusteringP2P (Muennighoff et al., 2023), designed for MTEB, comprises 10 splits of posts from StackExchange, each containing 5,000 to 10,000 entries, clustered by subreddit based on the combined content of titles and posts.
C.3 Reranking
MindSmall (Wu et al., 2020) dataset is a comprehensive English resource designed for research in news recommendation, focusing on ranking news articles based on the title of a currently read article to suggest related content.
SciDocsRR (Cohan et al., 2020) is a dataset designed for ranking related scientific papers using their titles as the primary basis for assessment.
StackOverflowDupQuestions (Liu et al., 2018) dataset focuses on identifying whether questions tagged with Java, JavaScript, and Python on Stack Overflow are duplicates of existing queries.
C.4 Retrieval
In retrieval task, each dataset consists of a corpus, queries, and a mapping of each query to relevant documents. The task goal is to find these relevant documents based on a given query. When evaluating, we first use the provided model to embed queries and corpus documents and then calculate the cosine similarity to obtain relevance scores and rank the corpus documents for each query based on these scores. The evaluation metrics consists of nDCG@k, MRR@k, MAP@k, precision@k, and recall@k, with nDCG@10 as the primary metric.
C.4.1 Beir Retrival
NFCorpus (Boteva et al., 2016) is a dataset that includes natural language queries sourced from NutritionFacts, paired with annotated medical documents from PubMed, utilizing the original splits from various types of content from NF, such as videos, blogs, and Q&A posts.
SciFact (Wadden et al., 2020) dataset evaluates scientific claims by matching them with evidence sourced from research literature, specifically utilizing a set of 300 test queries and the complete document collection from the original dataset.
C.4.2 LongEmbd Retrival
LongEmbed (Zhu et al., 2024) includes 4 real-world retrieval tasks curated from long-form QA and summarization. The document in LongEmbed is much longer compared to BEIR. Thus, it can effectively evaluate the capability of the embedding model on long texts.
LEMBNarrativeQARetrieval (Kočiskỳ et al., 2018) is a question-answering dataset featuring lengthy narratives, averaging 50,474 words, that challenge models to comprehend and extract information about characters and events dispersed throughout the stories.
LEMBQMSumRetrieval (Zhong et al., 2021) dataset focuses on generating summaries of meetings based on specific queries, necessitating the extraction and synthesis of relevant information from various segments of the conversation that cover multiple topics and participants.
LEMBSummScreenFDRetrieval (Chen et al., 2021) dataset consists of pairs of transcripts from TV series and their corresponding human-crafted summaries, requiring the integration of dispersed plot elements into concise narrative descriptions.
LEMBWikimQARetrieval (Ho et al., 2020) dataset is a complex question-answering resource that includes questions requiring up to five reasoning steps, designed using specific templates to encourage deep understanding rather than simple retrieval of information.
C.5 Semantic Textual Similarity (STS)
The task goal is to determine the similarity between a pair of sentences, where continuous scores serve as labels, with higher values indicating greater similarity. The provided model embeds the sentences, and their similarity is calculated using cosine similarity. The primary evaluation metric is the Spearman correlation (Reimers et al., 2016).
STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark (Agirre et al., 2012; 2013; Bandhakavi et al., 2014; Biçici, 2015; Nakov et al., 2016)  [^5] https://alt.qcri.org/semeval2017/task1/  [^6] https://competitions.codalab.org/competitions/33835  [^7] https://github.com/PhilipMay/stsb-multi-mt/ are collections of sentence pairs designed to evaluate semantic textual similarity, with the former set focused on monolingual English pairs and the latter two incorporating cross-lingual comparisons across multiple languages.
BIOSSES (Soğancıoğlu et al., 2017) comprises 100 pairs of sentences specifically focused on the biomedical domain.
SICK-R (Dadas et al., 2020) , which stands for Sentences Involving Compositional Knowledge, comprises 100,000 diverse sentence pairs that exhibit rich lexical, syntactic, and semantic characteristics.
C.6 Summarization
The input consists of a set of summaries written by humans and machines. The goal is to score the machine-generated summaries. Use the provided model to embed the summaries. Calculate the distance between each machine summary and all human summary embeddings. Retain similar scores as the model score for each individual machine-generated summary. Calculate the Spearman correlation based on cosine similarity (Reimers et al., 2016) as the main metric.
SummEval (Fabbri et al., 2021) consists of summaries produced by advanced summarization models trained on CNN and DailyMail articles.
Table 4: Statistics of the experimental datasets used in the work.
@@ C.6 Summarization
Appendix D Embedding Models
ANCE (Xiong et al., 2021) enhances dense retrieval by selecting challenging negative samples from the entire corpus and asynchronously updating the Approximate Nearest Neighbor (ANN) index with each training iteration, using a context window size of 512.
GTR (Ni et al., 2022) improves dual encoder performance for retrieval tasks by scaling up model size while keeping a fixed bottleneck embedding, leading to significant improvements in out-of-domain generalization, all within a context window size of 512.
GIST (Solatorio, 2024) consistently improves performance across different model sizes by leveraging the strengths of large, resource-intensive models to enhance smaller ones, making advanced AI technologies more accessible and cost-effective, all within a context window size of 512.
BGE (Xiao et al., 2023) offers a range of well-trained embedding models based on a BERT-like architecture, enabling users to balance performance and efficiency for various applications while also allowing easy fine-tuning. In our experiments, we use the gte-base-en-v1.5 model, which operates with a context window size of 512.
E5 (Zhu et al., 2024) is a long-context embedding model fine-tuned to support 4k token inputs while maintaining the original performance for shorter contexts, designed to advance research in long-context embedding technologies. It uses a context window size of 4k.
Table 5: Embedding models used in the experiments.

### v2

@@ B.3 Proof of Corollary 4
where $\alpha_{1}$ and $\alpha_{2}$ represent the maximum values in the frequency domain of $\mathcal{HC}\left[{\bm{x}_{1}}\right]$ and $\mathcal{HC}\left[{\bm{x}_{2}}\right]$ and $\alpha$ represent the value of $\mathcal{DC}\left[{\bm{x}_{1}}\right]$ and $\mathcal{DC}\left[{\bm{x}_{2}}\right]$, respectively, after applying the discrete Fourier transform. Eqn. 24 leverages that $\mathcal{HC}\left[{\cdot}\right]$ and $\mathcal{DC}\left[{\cdot}\right]$ are orthogonal. Eqn. 25 leverages the assumption that the mean of word embeddings in natural language texts maintains a relatively consistent representation. Thus $\mathcal{HC}\left[{\bm{x}_{1}}\right]$ and $\mathcal{HC}\left[{\bm{x}_{2}}\right]$ have the same value $\alpha$ after applying the discrete Fourier transform. Finally, according to Theorem 3, $\alpha_{1}$ and $\alpha_{2}$ will gradually decrease with $n$ grows, leading to a higher cosine similarity between $\bm{x}_{1}$ and $\bm{x}_{2}$. ∎
Appendix C More Experiments and Analysis
C.1 Rewriting Process
To investigate the differences in embedding distributions across texts of varying lengths, we use the Llama3 (i.e., Llama-3.1-8B-InstructDubey et al. (2024) model to rewrite the texts. Specifically, we used two prompts, “Please express the given text in one sentence. No more than 10 tokens. {{original text}}” and “Please use few sentences to summarize the given text. {{original text}}”, to summarize the texts. This rewriting allows the texts to retain the same semantics while having shorter lengths. By studying the differences in texts of varying lengths, we conclude that the cause of the Length Collapse phenomenon is that longer texts cluster to each other in embedding space.
C.2 Details on Figure 6
To verify Theorem 2, we illustrate the high-frequency intensity of each layer’s output along with its theoretical upper limit. Our visualization is based on the official checkpoint of 12-layer ANCE, GIST, and BGE. We use a logarithmic scale for the purpose of a better view. Let $\bm{X}_{l}$ denote the output of the $l$-th layer. For red line, we directly calculate $\log(\lVert\mathcal{HC}\left[{\bm{X}_{l+1}}\right]\rVert_{F}/\lVert\mathcal{HC}\left[{\bm{X}_{l}}\right]\rVert_{F})$ at each layer. In practice, models typically employ multi-head attention, so we replace $\lVert\bm{W}_{V}\rVert_{2}$ in Eqn. 1 with $\sigma_{1}H$, where $\sigma_{1}=\max_{h=1}^{H}\lVert\bm{W}_{V}^{h}\rVert_{2}$. For the blue line, we obtain the coefficient $\sigma_{1}$ with respect to network parameters and apply the logarithmic scale. To summarize, Figure 6 implies a convergence rate, which is consistent with our Theorem 2.
Figure 6: Visualization of the intensity of high-frequency components and their theoretical upper bounds. The blue line is defined by $\log(\lVert\mathcal{HC}\left[{\bm{X}_{l+1}}\right]\rVert_{F}/\lVert\mathcal{HC}\left[{\bm{X}_{l}}\right]\rVert_{F})$, and the red line is estimated using the results in Theorem 2.
C.3 Details on Figure 7
To verify Theorem 3, we visualize the value of $\sigma_{a}$ across different text lengths. Our visualization is based on the texts from NFCorpus. We sample 100 samples for each bin from 0 to 500 with a bin size of 50. The $\sigma_{a}$ value is computed as the average of the $\sigma_{a}$ based on the attention of all heads before the output of the final layer. To summarize, Figure 7 implies that $\sigma_{a}$ shows a decreasing trend as $n$ increases, which is consistent with our Theorem 3. Moreover, $\sigma_{a}$ also increases as $\tau$ decreases, further validating our proposed method, TempScale. E5 model shows an increasing $\sigma_{a}$ with length, which may be due to anomalies in the deep layer attention patterns.
(a) ANCE
(b) GTR
(c) GSIT
(d) BGE
(e) E5
C.4 More Analysis about Assumption in Theorem 4
Figure 8: The cosine similarity of embeddings of texts generated by repeating words “dog” and “cat”.
In Corollary 4, we hypothesize and verify that all-natural language sequences tend to have a relatively consistent representation. As a result, different texts tend to exhibit consistent low-pass signals after losing high-frequency information, leading to ultimately consistent text embeddings. This leads to an increase in cosine similarity for longer texts. However, as shown in Figure 8, we repeat the words “dog” and “word” $n$ times and calculate the similarity between these two texts. The results show that even when the sequences do not overlap, the text embeddings tend to converge to similar representations as the sequence length increases. This further demonstrates that Length Collapse causes completely different sequences to converge toward similarity.
C.5 More Discussion about Other Works
Other Components in Transformer. After discussing how self-attention contributes to Length Collapse, we proceed to examine the influence of other modules in the transformer, such as multi-head, residual, and FFN. Fortunately, previous work Wang et al. (2022b) has talked about whether these components can effectively alleviate the low-pass filtering drawbacks. The proof demonstrates that while these components help preserve high-frequency signals, they do not alter the fact that the MSA block, as a whole, functions solely with the representational power of a low-pass filter. Furthermore, the ability of these models to preserve high-frequency signals is solely determined by their internal parameters and architecture, independent of the input text length. As a result, these modules do not impact our analysis of Length Collapse in practical models.
Difference from Over-Smoothing in Deeper Layers. In previous research Wang et al. (2022b), it has been noted that a self-attention module acts as a low-pass filter, causing input feature maps to gradually lose high-frequency signals as the model layers go deeper. Furthermore, other studies Oono and Suzuki (2020); Cai and Wang (2020) indicate that the node features of Graph Convolutional Networks (GCNs) can become exponentially trapped in the null space of the graph Laplacian matrix. The root cause of this phenomenon is that both graph Laplacian matrices and self-attention matrices consistently exhibit a dominant eigenvector, commonly referred to as the DC component. While these studies address over-smoothing in deeper layers, we focus on how the low-pass filtering process changes as the input sequence lengthens, specifically examining over-smoothing in longer sequences.
C.6 More discussions on Classification and Clustering Tasks
For data grouping tasks like classification and clustering, in the case of $N$-class tasks, we can think of the model as learning $N$ classification boundaries. The farther the text embedding is from the boundary, the closer the output probability approaches 1 or 0. As shown in Figure 9 (left), we plot the entropy of the model output probabilities across different length intervals. The model outputs higher entropy for longer texts, which may be because the embeddings of longer texts are positioned closer to the center of the space as described in Figure 1(b), resulting in a shorter distance to various classification boundaries. Meanwhile, in Figure 9 (right), we can also observe that accuracy and entropy follow the same trend: the model achieves higher accuracy when it has lower entropy. In other words, the model performs better if the text embeddings are farther from the boundary. After applying TempScale, a decreased entropy will result in increased accuracy. This further supports the relationship between entropy and accuracy in classification tasks. Moreover, if the model exhibits a more severe Length Collapse phenomenon, meaning a greater performance drop on longer texts, the more performance improvement it experiences after applying TempScale. This suggests that one possible reason TempScale is effective on shorter text datasets is that it adjusts the distribution differences between texts of varying lengths. As shown in Figure 1(c), after applying TempScale, the distribution of text embeddings across different length intervals becomes more uniform, which facilitates the model in learning length-agnostic parameters.
C.7 More Results on Longer Texts
Figure 10: Results about performance difference between long query and short query across varying temperature $\tau$ on SummScreenFD dataset.
Can the temperature be set based on the length of the text? In the previous experiments, we use the same temperature $\tau$ for scaling all texts in the same task. However, our analysis indicates that texts of different lengths have varying filtering rates, so a natural idea is to use different temperatures for texts of different lengths. As shown in Figure 12, we plot the performance trend for texts under the same settings in Figure 9 as the temperature varies. The results indicate that a higher temperature is optimal for short texts, while a lower temperature is preferable for long texts, as confirmed by the results in Table 2. When performing retrieval tasks, we can also set different temperatures for queries and documents to achieve better performance. As shown in Figure 11, on the QMSum dataset, we can consistently achieve better performance by setting a lower temperature for queries. Moreover, as shown in Figure 13, compared to the QMSum dataset, SummScreenFD requires a lower temperature for scaling the document due to its longer length. This further supports the conclusion that longer texts require a lower temperature for scaling. In Figure 10, we observe a similar phenomenon across other models. Except for ANCE, the performance of other models on long queries decreases as the temperature decreases. This suggests that long queries require a lower temperature to mitigate the Length Collapse phenomenon. However, ANCE’s performance degradation with decreasing temperature is likely attributable to its inherent limitations in processing long texts.
Figure 12: Results about performance difference between long query and short query across varying temperature $\tau$ on AmazonPolarity dataset.
Figure 13: Relative performance compared to the raw results with varying query (Q Temperature) and document (D Temperature) temperatures using different models on SummScreenFD
Appendix D Length Collapse in LLM-based Embedding Models
To explore whether Length Collapse also occurs in long-context LLMs, we select three widely used LLM-based embedding models from the MTEB benchmark—bge-multilingual-gemma2 Chen et al. (2024a), NV-Embed-v2 Lee et al. (2024), and e5-mistral-7b-instruct Wang et al. (2022a)—all of which rank within the top 20 on the MTEB leaderboard. We conduct experiments using three commonly used long-text datasets, including nq and hotpotqa in LongRAG Jiang et al. (2024) and LongAlpaca-12k Chen et al. (2024b), which are selected based on relevance to the keyword "long" and chosen for their evenly distributed text lengths. Specifically, we analyze shifts in embedding space across different text length intervals by calculating the average Euclidean distance of each embedding from the central embedding (the mean of all embeddings) and computing cosine similarity between each pair of embeddings as in Table 4 and Table 5.
Dataset Model 0-1000 1000-2000 2000-3000
nq bge-multilingual-gemma2 0.94 0.92 0.91
NV-Embed-v2 0.98 0.91 0.89
e5-mistral-7b-instruct 0.68 0.64 0.64
hotpotqa bge-multilingual-gemma2 0.93 0.88 0.87
NV-Embed-v2 0.95 0.84 0.79
e5-mistral-7b-instruct 0.70 0.61 0.57
LongAlpaca-12k bge-multilingual-gemma2 0.71 0.48 0.49
NV-Embed-v2 0.89 0.86 0.87
e5-mistral-7b-instruct 0.69 0.53 0.53
Table 4: Euclidean distance results for different datasets and LLM-based embedding models.
Dataset Model 0-1000 1000-2000 2000-3000
nq bge-multilingual-gemma2 0.39 0.42 0.45
NV-Embed-v2 0.32 0.47 0.59
e5-mistral-7b-instruct 0.75 0.77 0.78
hotpotqa bge-multilingual-gemma2 0.51 0.52 0.59
NV-Embed-v2 0.48 0.59 0.69
e5-mistral-7b-instruct 0.75 0.81 0.85
LongAlpaca-12k bge-multilingual-gemma2 0.76 0.90 0.89
NV-Embed-v2 0.48 0.65 0.68
e5-mistral-7b-instruct 0.78 0.87 0.87
Table 5: Pair cosine similarities results for different datasets and LLM-based embedding models.
The experimental results show that as text length increases, the embeddings from different LLM-based embedding models exhibit a gradual convergence trend, with the average Euclidean distance between embeddings decreasing and pairwise cosine similarity increasing. This indicates that even mainstream long-context LLM embedding models tend to experience embedding convergence (Length Collapse) and reduced distinctiveness in long-text processing due to the low-pass filtering effect.
Appendix E More Discussions about TempScale
E.1 Relationship with Contrastive Learning
Background and Motivation. Contrastive learning is widely used to address embedding space anisotropy, reducing high similarity between random text samples of varying lengths by maximizing distances among negative pairs and aligning positive pairs. While it improves embedding quality in many applications, its impact on Length Collapse remains unexplored. This section examines how contrastive learning affects Length Collapse, assessing its contributions and limitations.
As shown in previous works Wang and Isola (2020), the InfoNCE loss, when scaled to a large number of negative samples, can be decomposed into two primary components: Alignment and Uniformity. Alignment ensures that positive pairs—texts with similar content—are close in the embedding space, while Uniformity spreads embeddings of negative pairs to prevent them from clustering excessively. Mathematically, as the number of negative samples $M\to\infty$, the normalized InfoNCE loss can be expressed as:
$\lim_{M\rightarrow\infty}L(f,\tau)-\log M$ $=-\frac{1}{\tau}E_{\left(x,x^{+}\right)\sim p_{\text{pos}}}\left[f(x)^{T}f\left(x^{+}\right)\right]$
$+E_{x\sim p_{\text{data}}}\left[\log E_{x^{-}\sim p_{\text{data}}^{-}}\left[e^{f(x)^{T}f\left(x^{-}\right)/\tau}\right]\right].$
These properties make contrastive learning especially effective in tasks such as retrieval, where maximizing inter-sample variance is crucial.
E.1.1 Why Contrastive Learning Cannot Fully Address Length Collapse
While contrastive learning alleviates high similarity issues across all text lengths, it does not entirely resolve the Length Collapse issue inherent in PLM-based models. Length Collapse arises from self-attention’s tendency to push embeddings for longer texts toward a concentrated representation space. This characteristic is unaffected by contrastive learning’s alignment or uniformity mechanisms because contrastive learning optimizes the relative positioning of positive and negative pairs rather than mitigating the length-induced clustering trend.
E.2 Comparison with Other Post-processing Techniques
In addition to TempScale, several post-processing methods have been proposed to address high similarity in embeddings, such as the Flow Function Li et al. (2020) and Whitening Su et al. (2021). For comparative analysis, we implemented the last2avg version of the Flow Function, as the full flow version requires additional training, and the Whitening method is based on its original formulation. The results, summarized in Table 6, highlight that these methods do not perform as effectively in our specific application. These results indicate that Flow and Whitening, originally developed for standard BERT embeddings, are less effective with the fine-tuned Transformer model used in this paper, which includes additional normalization layers. While TempScale reduces similarity in long-sequence embeddings, achieving distributional consistency across different sequence lengths is more critical for performance. Therefore, lowering similarity alone may not significantly improve downstream tasks.
Model Rerank. Summ. Class. Clust. LongEmbdRetr. STS BeirRetr. Avg.
ANCE 49.09 29.58 55.27 33.04 34.02 66.32 36.87 43.45
+Flow 48.49 29.57 56.34 32.11 31.92 66.02 36.57 43.00
+Whitening 25.16 21.76 25.82 3.21 1.29 2.78 0.79 11.54
GTR 54.23 29.67 55.10 38.65 37.33 70.11 44.98 47.15
+Flow 37.49 28.09 41.98 21.87 7.65 35.55 1.35 24.85
+Whitening 24.63 20.69 25.71 3.06 1.35 -3.02 0.63 10.44
GIST 58.55 31.14 64.75 44.77 38.21 75.61 52.77 52.26
+Flow 58.16 30.56 65.68 44.67 35.83 74.98 51.40 51.61
+Whitening 24.08 24.26 20.06 9.83 1.78 -4.31 0.81 10.93
BGE 58.87 31.03 64.79 45.80 37.46 75.88 55.29 52.73
+Flow 58.48 30.73 64.77 45.28 36.95 74.43 54.52 52.16
+Whitening 24.75 19.61 16.50 4.23 2.39 1.60 0.47 9.94
E5-4K 53.12 30.58 61.72 41.01 56.01 71.77 47.22 51.63
+Flow 52.48 29.82 62.33 39.95 44.71 68.55 30.68 46.93
+Whitening 24.94 22.14 22.73 4.04 1.85 1.96 0.56 11.17
Table 6: Performance comparison of different post-processing methods across various tasks
E.3 Comparison with Other Similar Long-Text Methods
To thoroughly evaluate our TempScale approach, we compare it with two related attention scaling methods: $\text{softmax}\left(\frac{\text{logn}}{\tau\sqrt{d}}QK^{T}\right)V$ (Method1) Chiang and Cholak (2022) and YaRN Peng et al. (2024a). In YaRN, the scaling formula for the attention matrix is given as $\text{softmax}\left(\frac{1}{\tau\sqrt{d}}QK^{T}\right)V$, where $\frac{1}{\tau}=0.1\ln{s}+1$, and $s=\frac{L^{\prime}}{L}$, with $L^{\prime}$ representing the extended context window length and $L$ the original context window length. These methods propose different scaling strategies, but they differ in their theoretical motivations, applications, and experimental outcomes.
While Method1, YaRN, and TempScale share a similar structural approach, they tackle different challenges. Method1 is for binary classification tasks, specifically determining if the first string in a sequence of binary strings is a “1”. It focuses on resolving straightforward decision-making problems in binary data. YaRN, by contrast, extends the context window of LLMs without requiring retraining, particularly adjusting the attention mechanism to handle larger contexts. TempScale addresses Length Collapse in embedding models for long texts, improving performance without retraining by preserving distinct representations as text length increases. Moreover, Method1 offers a solution using a specially designed Transformer example, but its motivation is limited by the specificity of this example. In contrast, YaRN introduces an empirically derived scaling formula. TempScale is based on a rigorous low-pass filtering analysis, giving it a strong theoretical foundation and an intuitive explanation.
Model Rerank. Summ. Class. Clust. LongEmbdRetr. STS BeirRetr. Avg.
ANCE 49.09 29.58 55.27 33.04 34.02 66.32 36.87 43.45
+Method1 41.80 29.21 48.21 20.72 6.23 53.40 7.74 29.61
+YaRN($\lambda=0.0001$) 49.09 29.58 55.62 33.01 34.02 66.32 36.86 43.50
+YaRN($\lambda=0.001$) 49.08 29.57 55.62 33.05 33.96 66.32 36.87 43.49
+YaRN($\lambda=0.01$) 49.10 29.31 55.65 32.95 33.92 66.27 36.87 43.44
+YaRN($\lambda=0.1$) 48.95 29.30 55.45 32.65 32.31 65.72 35.77 42.88
+YaRN($\lambda=1$) 45.71 29.28 52.80 26.16 10.59 60.15 13.34 34.00
GTR 54.23 29.67 55.10 38.65 37.33 70.11 44.98 47.15
+Method1 38.12 28.44 40.34 14.46 3.61 47.09 0.82 24.70
+YaRN($\lambda=0.0001$) 54.23 29.68 55.06 38.30 37.42 70.11 44.98 47.11
+YaRN($\lambda=0.001$) 54.23 29.71 55.08 38.29 37.56 70.11 44.97 47.13
+YaRN($\lambda=0.01$) 54.23 29.71 55.04 38.63 37.39 70.06 45.00 47.15
+YaRN($\lambda=0.1$) 54.12 29.34 54.86 34.07 36.53 69.86 43.06 45.98
+YaRN($\lambda=1$) 55.19 29.00 47.27 19.06 3.99 61.87 1.46 31.12
GIST 58.55 31.14 64.75 44.77 38.21 75.61 52.77 52.26
+Method1 58.64 28.26 51.05 28.11 5.42 62.43 7.51 34.49
+YaRN($\lambda=0.0001$) 58.55 31.14 64.26 44.75 38.21 75.61 52.78 52.19
+YaRN($\lambda=0.001$) 58.55 31.12 64.28 44.75 38.20 75.61 52.78 52.18
+YaRN($\lambda=0.01$) 58.53 31.26 64.24 44.78 38.05 75.60 52.79 52.18
+YaRN($\lambda=0.1$) 58.45 30.36 63.86 44.62 37.34 75.31 52.10 51.72
+YaRN($\lambda=1$) 55.75 26.76 58.79 36.95 11.85 69.00 23.59 40.38
BGE 58.87 31.03 64.79 45.80 37.46 75.88 55.29 52.73
+Method1 37.78 29.17 37.80 14.36 2.10 43.72 0.84 23.68
+YaRN($\lambda=0.0001$) 58.86 31.04 64.78 45.77 37.45 75.88 55.29 52.72
+YaRN($\lambda=0.001$) 58.86 30.96 64.78 45.75 37.47 75.87 55.28 52.71
+YaRN($\lambda=0.01$) 58.85 31.02 64.73 45.61 37.26 75.86 55.22 52.65
+YaRN($\lambda=0.1$) 58.86 30.90 64.51 45.19 36.55 75.55 54.96 52.36
+YaRN($\lambda=1$) 52.65 29.09 50.51 25.50 2.36 64.20 2.14 32.35
E5-4K 53.12 30.58 61.72 41.01 56.01 71.77 47.22 51.63
+Method1 40.90 24.11 42.85 13.64 3.24 41.62 1.05 23.92
+YaRN($\lambda=0.0001$) 53.12 30.57 61.78 40.77 56.01 71.77 47.22 51.61
+YaRN($\lambda=0.001$) 53.11 30.55 61.78 40.75 55.98 71.77 47.22 51.59
+YaRN($\lambda=0.01$) 53.07 30.42 61.73 40.61 55.53 71.71 47.21 51.47
+YaRN($\lambda=0.1$) 52.64 30.20 61.51 40.19 48.39 70.68 46.36 50.00
+YaRN($\lambda=1$) 45.00 29.01 50.81 17.99 2.91 57.27 1.46 29.21
Table 7: Average main metric on MTEB and LongEmbd across Method1 and YaRN.
To effectively compare the performance of the above methods, we tested them on the MTEB benchmark. Since we are not modifying the context window length, we adapt YaRN as $\text{softmax}\left(\frac{\tau}{\sqrt{d}}QK^{T}\right)V$, with $\tau=\lambda\log n+1$. We explore values for $\lambda$ in the range $\{0.0001,0.001,0.01,0.1,1\}$. This setup is reasonable, as our findings indicate that longer texts generally benefit from a smaller temperature scale. The experimental results are as shown in Table 7.
The experimental results indicate that neither Method1 nor YaRN effectively adapts to the embedding model scenario. (Although YaRN shows slight improvement when the $\lambda$ value is small, in this case, Method 2 degenerates into TempScale.) Some potential reasons are as follows: A plausible explanation is that while both methods perform finer scaling on the attention matrix, applying different temperature adjustments across varying text lengths may lead to embeddings from different lengths falling into distinct distributions, which is unfavorable for downstream tasks. Moreover, in generation tasks, Method1 and YaRN succeed likely because of differences in output requirements. For these tasks, the model outputs a probability distribution and samples from it. Minor perturbations generally don’t affect the token output significantly; even if one sequence’s token distribution changes, it doesn’t impact the output of other sequences. In contrast, any substantial change in a single embedding for an embedding model can directly affect the entire downstream performance. For instance, in classification tasks, a classifier model relies on embeddings as input, and in retrieval tasks, a change in embedding impacts document ranking. Overall, for embedding tasks, simply applying a single temperature adjustment better maintains the overall embedding distribution, helping mitigate Length Collapse and achieve better results across various downstream applications.
Appendix F Datasets and Evaluation Metrics
@@
Table 8: Statistics of the experimental datasets used in the work.